OpenStudy (anonymous):

Drawing three queens in a row from a standard deck of cards when the drawn card is not returned to the deck each time..............The probability is________???? (type an integer or simplified fraction) does this make sense?? i need help

OpenStudy (anonymous):

\[\frac{4}{52}\times \frac{3}{51}\times \frac{2}{50}\] rather unlikely you can reduce that a bit

OpenStudy (anonymous):

thanks maybe i didnt word it right?? i dont know but the answer turned out to be 1/5525??? thats what it said anyway